A) A quality analyst wants to examine a packaging process that has an overall average weight of 849 lb. The average range is 10

B) A quality analyst wants to examine a packaging process that has an overall average weight of 849 lb. The average range based
Answer #1

(b)

Average range (R-bar) = 14 lb

For sample size of 12, the factor for control limits for range(D3) = 0.283 (obtained from the table of constants for x-bar and R chart)

LCL for R-chart = D3 × R-bar = 0.283 × 14 = 3.962